OpenStudy (anonymous):

Solve |-3x-1| (2) -2x-5(0)

OpenStudy (anonymous):

@Cryxz Fix the typo in your post!

OpenStudy (anonymous):

there is no typo. the question is # 8.

OpenStudy (anonymous):

Given f(x) = 2x - 5, g(x) = | -3x -1| -> f(0) = -5 -> g (2) = | -6 - 1| = 7 => g (2) - f(0) = 7 + 5 = 12

OpenStudy (anonymous):

@Cryxz I consider your post is typo, because no way I can understand it without the attachment :/

OpenStudy (anonymous):

left to right or is it up to down?

OpenStudy (anonymous):

That's 2 separated functions!

OpenStudy (anonymous):

Okay. Thanks so much! :)